always have issues with fractions have problems with this equation. solve by substitution. 3x-2y=3 x-1/3y=8

Answer by Alan3354(69443)

3x-2y=3
x-1/3y=8 times 3 --> 3x - y = 24 **** No fractions.
-----------
3x - 2y = 3
3x - y = 24